# VISTA: Knowledge-Driven Interpretable Vessel Trajectory Imputation via Large Language Models

### v1

@@ 6. Conclusion
We present VISTA, a knowledge driven and interpretable vessel trajectory imputation framework that returns reconstructed trajectory segments together with supporting knowledge for downstream use. VISTA defines this support as underlying knowledge that combines Structured Data-derived Knowledge (SDK) from AIS data with the implicit knowledge encoded in large language models (LLMs). VISTA organizes the SDK as a Structured Data-derived Knowledge Graph (SD-KG) and applies it in a data-knowledge-data loop. When going from data to knowledge, VISTA distills vessel attributes, behavior patterns, and validated imputation functions from raw AIS data and inserts them into the SD-KG. When going from knowledge to data, it uses this structured knowledge to estimate behavior, select and execute imputation methods, and generate human readable explanations grounded in maritime rules and operations. In addition, to ensure efficient processing of large scale AIS data, a workflow manager layer enables parallel scheduling, anomaly guards, and de-duplication, which reduces LLM latency and stabilizes throughput. Experiments on two AIS datasets show state-of-the-art accuracy and substantial efficiency gains, and the produced knowledge cues enhance downstream analysis; ablations show that the workflow management layer effectively reduces runtime through parallel scheduling and preserves the compactness of the SD-KG through the de-redundancy strategy.
In future work we plan to extend VISTA into a unified maritime trajectory management framework that supports multiple data management tasks, including trajectory imputation, anomaly detection, and trajectory prediction.
7. Acknowledgment
The research leading to the results presented in this paper has received funding from the European Union’s funded Project MobiSpaces under grant agreement no 101070279.
@@ A.1. The SD-KG Construction and Maintenance
A.1.1. Behavior Patterns Extraction
Figure 8 presents Behavior Abstraction Prompt, which operationalizes the behavior abstraction in Section 4.2.2 by mapping raw motion signals into a discrete, human-interpretable pattern for downstream imputation. The kinematic tokens $(p^{s},p^{\theta},p^{\psi})$ are instantiated as speed_pattern, course_pattern, and heading_pattern, respectively. The prompt explicitly instructs the LLM to analyze the temporal evolution within the provided {trajectory_data} and to choose from finite vocabularies ({speed_dict}, {course_dict}, {heading_dict}) before creating any new entries, thereby constraining vocabulary growth and enabling deduplication of semantically equivalent labels. And the intent token $p^{i}$ is realized through two complementary fields: a high-level intent (e.g., ‘navigating’) and a context-grounded navigation intent (e.g., ‘the vessel is maintaining its course’). Both are inferred from static and contextual cues, navigation status $\eta$, vessel type $\kappa$, and cargo type $\chi$ and spatial context $\sigma$ (obtained from Static and Spatial Encoder) in trajectory_data; a finite set {intent_dict} regulates label usage to preserve interpretability and retrieval efficiency. Finally, the output schema mandates a single structured pattern enclosed in triple quotes, ensuring robust parsing and consistent downstream indexing.
Imputation Method Builder [TASK] You are an expert in maritime trajectory analysis and spatial-temporal modeling, specialized in developing interpretable algorithms for vessel movement reconstruction. You are given vessel trajectory data and are asked to generate a spatial_function that estimates missing latitude and longitude positions based on known trajectory features and motion patterns. You need to adhere requirements: The spatial_function should be a single-line, directly executable Python function. The spatial_function is encouraged to choose from a variety of path models, not just linear interpolation. The spatial_function must compute a sequence of intermediate points using the provided start, end and Time_interval. start: A tuple representing the geographic coordinates (latitude, longitude) immediately before the missing data block.
@@ A.1.3. Overall Procedure of SD-KG Construction and Maintenance
A.2. Knowledge Driven Trajectory Imputation
A.2.1. Behavior Pattern Selection and Rationale.
Figure 10 presents the prompt used for LLM-driven Behavior Pattern Selection and Rationale described in Section 4.3.1. This prompt guides the LLM to identify the most behavior pattern for a missing trajectory segment and to articulate the reasoning behind this choice.
@@ A.2. Knowledge Driven Trajectory Imputation
Method Selector and Rationale [TASK] You are an expert in maritime spatial-temporal modeling and trajectory reconstruction, responsible for evaluating and selecting the most suitable spatial function for accurate AIS trajectory imputation. Please select the most suitable spatial function for imputing missing latitude and longitude.
@@
A.2.2. Imputation Method Selection and Rationale.
@@
A strict output schema enforces machine-readability and interpretability: the model must return a triple-quoted block containing 1) Selected Function ID, 2) Statistical Support, which reports only statistical evidence (translated from association frequencies into probabilities) without mentioning graph structures or edge weights explicitly, each probability is accompanied by its calculation from weight proportions to prevent unverifiable claims—and and 3) Reasoning, a technical justification of why the method matches the kinematic requirements of $v_{b}^{*}$. This schema ensures the selected method $v_{f}^{*}\in\mathcal{C}_{f}^{K}$ is both verifiable (via explicit probability construction from $\mathcal{G}_{\iota,f}^{K}$) and actionable (via the prescribed imputation procedure), while the resulting rationale is stored in $\mathcal{J}_{\iota}^{k,f}$ for downstream auditing and integration.
@@
A.2.3. Underlying Case for Human-friendly Explanation.
@@ 33 $\mathcal{R}\leftarrow\mathcal{R}\cup\{R_{\iota}\}$
A.2.4. Overall Procedure of Knowledge Driven Trajectory Imputation
Algorithm 1 summarizes the dataset-level inference workflow that imputes all gaps using the SD-KG as a knowledge substrate; the concrete pseudocode is given in Algorithm 2. The procedure iterates over vessels and their segmented streams, retrieves boundary knowledge for each gap, selects a behavior and a method with graph-supported rationales, executes the method to reconstruct the missing segment, and composes a human-friendly explanation.
@@ A.3. Workflow Management Layer
A.3.1. The de-reduplication strategy of De-redundancy Processor
@@ A.3.2. Overall Procedure of SD-KG Construction Workflow Manager
Algorithm 3 orchestrates knowledge extraction with a two–stage parallel workflow (extraction $\rightarrow$ de-redundancy) connected by stack-based scheduling and batch barriers. Lines 1–6 initialize the workflow by constructing two stacks: the compute stack $\mathcal{S}_{c}$ and the de-duplication stack $\mathcal{S}_{d}$. All minimal segments $\mathbf{S}_{\iota}^{k}$ are pushed to $\mathcal{S}_{c}$ in timestamp order, each as a job tuple $(\iota,k,\mathbf{S}_{\iota}^{k},c)$ with an initial retry counter c=0c{=}0. Lines 7–12 launch batched parallel extraction. While $\mathcal{S}_{c}$ is non-empty, a batch of at most $b$ jobs is popped and executed concurrently (ParFor), where each job calls ExtractKU to produce a knowledge unit $u_{\iota}^{k}$ for segment $\mathbf{S}_{\iota}^{k}$. A synchronization barrier follows each batch to ensure consistent collection of results. Lines 13–21 perform anomaly detection and bounded retry. Each extracted unit is examined by Anomaly_Detect. If an anomaly is found and the retry counter satisfies $c<\varrho_{c}$, the job is re-queued to $\mathcal{S}_{c}$ with counter incremented to c+1c{+}1. Otherwise, the unit is admitted into the valid batch $\mathcal{B}_{u}$. Jobs that exceed $\varrho_{c}$ retries are quarantined. The validated batch $\mathcal{B}_{u}$ is then pushed onto $\mathcal{S}_{d}$ for de-redundancy. Lines 22–29 execute the parallel de-redundancy stage. Each batch popped from $\mathcal{S}_{d}$ is processed in parallel by Deredund_Processor, which merges semantically equivalent behavior tokens and function implementations to produce de-duplicated results $\mathcal{B}_{u}^{d}$. All cleaned units are accumulated into $\mathcal{U}_{d}$, enabling scalable consolidation independent of extraction. Lines 30–33 commit the consolidated results to the SD-KG. Each $(v_{s},v_{b},v_{f})\in\mathcal{U}_{d}$ is inserted or reused within $\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}$, and the corresponding edges $(v_{s},v_{b})$ and $(v_{b},v_{f})$ are updated via EdgeInc. Each per-vessel sequence $\mathcal{U}_{d,\iota}$ is maintained in temporal order to align with segment index $k$. Line 34 returns the updated SD-KG $\mathcal{G}_{d}$ and the complete set of de-duplicated knowledge units $\mathcal{U}_{d}$.
@@ B.1. Datasets
We use two AIS datasets: AIS-DK from the Danish Maritime Authority and AIS-US from the National Oceanic and Atmospheric Administration. AIS-DK covers Danish waters in March 2024, including major routes in the Baltic and North Seas; it contains 10,000 vessel sequences and 2,000,000 records, with an average sequence duration of 0.5 hours across 348 vessels. AIS-US covers U.S. coastal waters in April 2024 with dense traffic near major ports; it contains 10,000 vessel and 2,000,000 records, with an average sequence duration of 2.8 hours across 4,723 vessels. Both datasets include diverse vessel types (e.g., Cargo, Tanker, Passenger), providing representative coverage for evaluation. To construct these datasets, we first downloaded the raw AIS data from the official sources and performed data cleaning to remove abnormal or incomplete records. We then partitioned each vessel trajectory—March 2024 for AIS-DK and April 2024 for AIS-US—into fixed-length segments of 200 records, and uniformly sampled 10,000 segments to build the evaluation corpus.
@@ B.2. Hyperparameters and Implementations
Lin-ITP, Akima spline, and Kalman Filter are implemented in Python using pandas and pykalman. MH-GIN (Liu et al., 2025), Multi-task AIS (Nguyen et al., 2018), and KAMEL (Musleh and Mokbel, 2023) follow the authors’ default configurations. GLM-4.5, GLM 4.5-air, Qwen-Plus (snapshot 0112), and Qwen-flash (snapshot 2025-07-28) are accessed through the Aliyun Bailian API (Aliyun Teams, 2025) with thinking mode enabled.
@@ B.3. Evaluation Metrics
We use the Haversine geodesic distance as the joint spatial error and report it in kilometer. The formulas are:
@@ B.4. Baseline Methods
We compare VISTA with three classes of imputation functions that align with Table 1: rule based trajectory imputation, deep learning based trajectory imputation, and LLM based trajectory imputation.
@@
For rule based trajectory imputation, we include Lin-ITP (Widyantara et al., 2023), Akima Spline (Zaman et al., 2023), and Kalman Filter (Wang et al., 2023). 1) Lin-ITP linearly interpolates longitude and latitude between boundary observations, implicitly assuming constant velocity and heading, and is simple and fast. 2) Akima Spline constructs a piecewise cubic polynomial using local slopes estimated from neighboring points, avoiding oscillations and overshooting common in standard cubic splines while maintaining smooth curvature continuity. It performs better on irregular or nonuniform motion patterns. 3) Kalman Filter treats motion as a linear Gaussian state space with constant velocity dynamics; forward filtering with backward smoothing recovers latent states and positions and is most effective when kinematics are near linear under moderate Gaussian noise.
@@
For deep learning based trajectory imputation, we include MH-GIN (Liu et al., 2025) and Multi-task AIS (Nguyen et al., 2018). MH-GIN is a multi scale heterogeneous graph imputation network for AIS streams: it extracts multi scale temporal features per attribute while preserving heterogeneous update rates, then constructs a multi scale heterogeneous graph to capture cross attribute dependencies and propagates information to fill missing values. Multi-task AIS is a recurrent framework with latent variable modeling and an embedding of AIS messages designed for high volume, noisy, and irregularly sampled data; it jointly supports trajectory reconstruction, anomaly detection, and vessel type identification.
@@
For LLM based trajectory imputation, we include KAMEL (Musleh and Mokbel, 2023), Qwen-plus-thinking, Qwen-flash-thinking, GLM-4.5-thinking, and GLM-4.5-air-thinking. KAMEL casts trajectory completion as a missing word problem with spatially aware tokenization and multi point masked infilling. Beyond KAMEL, to probe the out of the box capability of general LLMs on trajectory imputation without any task specific design, we adopt two state of the art families, Qwen (Qwen Team, 2025) and GLM 4.5 (Zeng et al., 2025). For each family we include a lightweight and a full capacity variant, and we enable thinking mode (Alibaba Cloud Team, 2025) throughout. This yields two paired comparisons under a unified setup: Qwen-plus-thinking vs Qwen-flash-thinking for capacity scaling, and GLM-4.5-air-thinking vs GLM-4.5-thinking for the trade off between latency and accuracy.
@@
B.5. Prompt of General Large Language Model for Vessel Trajectory Imputation.
@@
Figure 14 presents the baseline prompt used to elicit intermediate points for a missing trajectory segment from a general-purpose LLM. The prompt conditions the model on two boundary observations, previous point ${p}^{-}=[\lambda^{-},\phi^{-}]$ ({prev_str}) and next (first post-gap) point ${p}^{+}=[\lambda^{+},\phi^{+}]$ ({next_str}), and a target count $m$ of missing samples ({missing_length}). The model must output a strict JSON array of $m$ coordinates $\big([\lambda_{1},\phi_{1}],\ldots,[\lambda_{L},\phi_{m}]\big)$, each in [longitude, latitude] with six decimal places, without accompanying text. The constraints emphasize (i) smooth and continuous spatial progression consistent with vessel motion heuristics and (ii) temporal continuity for plausible interpolation between ${p}^{-}$ and ${p}^{+}$. The example block in Figure 14 specifies the exact output schema to ensure deterministic parsing and fair, method-agnostic evaluation of baseline LLMs.

### v2

@@ 6. Conclusion
We present VISTA, a knowledge-driven vessel trajectory imputation framework that equips repaired trajectories with structured repair provenance—queryable metadata recording what behavior pattern is estimated, how the segment is imputed, and why the behavior occurred. By grounding LLM reasoning in a data-verified Structured Data-derived Knowledge Graph (SD-KG) through a closed data–knowledge–data loop, VISTA ensures that behavior estimation, imputation method selection, and causal explanation generation are all anchored in empirically validated knowledge rather than unconstrained LLM conjecture. A workflow management layer coordinates the full pipeline with parallel scheduling, fault tolerance, and redundancy control, enabling efficient processing of large-scale AIS data. Experiments on two real-world AIS datasets show that VISTA outperforms the strongest baseline by 5%–91% in accuracy while reducing inference time by 51%–93%; ablation studies confirm that each LLM role contributes distinctively and that de-redundancy maintains a compact yet expressive SD-KG. In future work, we plan to extend VISTA into a unified maritime trajectory management framework supporting trajectory imputation, anomaly detection, and trajectory prediction.
7. Acknowledgments
This research was supported in part by the European Union-funded Project MobiSpaces under grant agreement no 101070279.
@@ A.1. The SD-KG Construction and Maintenance
A.1.1. Behavior Pattern Extraction
Figure 7 presents the Behavior Abstraction Prompt, which operationalizes the behavior abstraction in Section 4.2.2 by mapping raw motion signals into a discrete, human-interpretable pattern for downstream imputation. The kinematic tokens $(p^{s},p^{\theta},p^{\psi})$ are instantiated as speed_pattern, course_pattern, and heading_pattern, respectively. The prompt explicitly instructs the LLM to analyze the temporal evolution within the provided {trajectory_data} and to choose from finite vocabularies ({speed_dict}, {course_dict}, {heading_dict}) before creating any new entries, thereby constraining vocabulary growth and enabling deduplication of semantically equivalent labels. The intent token $p^{i}$ is realized through two complementary fields: a high-level intent (e.g., ‘navigating’) and a context-grounded navigation intent (e.g., ‘the vessel is maintaining its course’). Both are inferred from static and contextual cues, navigation status $\eta$, vessel type $\kappa$, cargo type $\chi$, and spatial context $\sigma$ (obtained from Static and Spatial Encoder) in trajectory_data; a finite set {intent_dict} regulates label usage to preserve interpretability and retrieval efficiency. Finally, the output schema mandates a single structured pattern enclosed in triple quotes, ensuring robust parsing and consistent downstream indexing.
Imputation Method Builder [TASK] You are an expert in maritime trajectory analysis and spatio-temporal modeling, specialized in developing interpretable algorithms for vessel movement reconstruction. You are given vessel trajectory data and are asked to generate a spatial_function that estimates missing latitude and longitude positions based on known trajectory features and motion patterns. You need to adhere requirements: The spatial_function should be a single-line, directly executable Python function. The spatial_function is encouraged to choose from a variety of path models, not just linear interpolation. The spatial_function must compute a sequence of intermediate points using the provided start, end and Time_interval. start: A tuple representing the geographic coordinates (latitude, longitude) immediately before the missing data block.
@@ A.1.3. Overall Procedure of SD-KG Construction and Maintenance
A.2. Knowledge-Driven Trajectory Imputation
A.2.1. Behavior Pattern Selection and Rationale
Figure 9 presents the prompt used for LLM-driven Behavior Pattern Selection and Rationale described in Section 4.3.1. This prompt guides the LLM to identify the most plausible behavior pattern for a missing trajectory segment and to articulate the reasoning behind this choice.
@@ A.2.1. Behavior Pattern Selection and Rationale
Method Selector and Rationale [TASK] You are an expert in maritime spatio-temporal modeling and trajectory reconstruction, responsible for evaluating and selecting the most suitable spatial function for accurate AIS trajectory imputation. Please select the most suitable spatial function for imputing missing latitude and longitude.
@@
A.2.2. Imputation Method Selection and Rationale
@@
A strict output schema enforces machine-readability and interpretability: the model must return a triple-quoted block containing 1) Selected Function ID, 2) Statistical Support, which reports only statistical evidence (translated from association frequencies into probabilities) without mentioning graph structures or edge weights explicitly, each probability is accompanied by its calculation from weight proportions to prevent unverifiable claims—and 3) Reasoning, a technical justification of why the method matches the kinematic requirements of $v_{b}^{*}$. This schema ensures the selected method $v_{f}^{*}\in\mathcal{C}_{f}^{K}$ is both verifiable (via explicit probability construction from $\mathcal{G}_{\iota,f}^{K}$) and actionable (via the prescribed imputation procedure), while the resulting rationale is stored in $\mathcal{J}_{\iota}^{k,f}$ for downstream auditing and integration.
@@
A.2.3. Explanation Composer for Human-Friendly Explanation
@@ 33 $\mathcal{R}\leftarrow\mathcal{R}\cup\{R_{\iota}\}$
A.2.4. Overall Procedure of Knowledge-Driven Trajectory Imputation
Algorithm 2 summarizes the dataset-level inference workflow that imputes all gaps using the SD-KG as a knowledge substrate. The procedure iterates over vessels and their segmented streams, retrieves boundary knowledge for each gap, selects a behavior and a method with graph-supported rationales, executes the method to reconstruct the missing segment, and composes a human-friendly explanation.
@@ A.3. Workflow Management Layer
A.3.1. De-redundancy Strategy
@@ A.3.2. Overall Procedure of SD-KG Construction Workflow Manager
Algorithm 3 orchestrates knowledge extraction with a two-stage parallel workflow (extraction $\rightarrow$ de-redundancy) connected by stack-based scheduling and batch barriers. Lines 1–6 initialize the workflow by constructing two stacks: the compute stack $\mathcal{S}_{c}$ and the de-duplication stack $\mathcal{S}_{d}$. All minimal segments $\mathbf{S}_{\iota}^{k}$ are pushed to $\mathcal{S}_{c}$ in timestamp order, each as a job tuple $(\iota,k,\mathbf{S}_{\iota}^{k},c)$ with an initial retry counter c=0c{=}0. Lines 7–12 launch batched parallel extraction. While $\mathcal{S}_{c}$ is non-empty, a batch of at most $b$ jobs is popped and executed concurrently (ParFor), where each job calls ExtractKU to produce a knowledge unit $u_{\iota}^{k}$ for segment $\mathbf{S}_{\iota}^{k}$. A synchronization barrier follows each batch to ensure consistent collection of results. Lines 13–21 perform anomaly detection and bounded retry. Each extracted unit is examined by Anomaly_Detect. If an anomaly is found and the retry counter satisfies $c<\varrho_{c}$, the job is re-queued to $\mathcal{S}_{c}$ with counter incremented to c+1c{+}1. Otherwise, the unit is admitted into the valid batch $\mathcal{B}_{u}$. Jobs that exceed $\varrho_{c}$ retries are quarantined. The validated batch $\mathcal{B}_{u}$ is then pushed onto $\mathcal{S}_{d}$ for de-redundancy. Lines 22–29 execute the parallel de-redundancy stage. Each batch popped from $\mathcal{S}_{d}$ is processed in parallel by Deredund_Processor, which merges semantically equivalent behavior tokens and function implementations to produce de-duplicated results $\mathcal{B}_{u}^{d}$. All cleaned units are accumulated into $\mathcal{U}_{d}$, enabling scalable consolidation independent of extraction. Lines 30–33 commit the consolidated results to the SD-KG. Each $(v_{s},v_{b},v_{f})\in\mathcal{U}_{d}$ is inserted or reused within $\mathcal{V}_{s},\mathcal{V}_{b},\mathcal{V}_{f}$, and the corresponding edges $(v_{s},v_{b})$ and $(v_{b},v_{f})$ are updated via EdgeInc. Each per-vessel sequence $\mathcal{U}_{d,\iota}$ is maintained in temporal order to align with segment index $k$. Line 34 returns the updated SD-KG $\mathcal{G}_{d}$ and the complete set of de-duplicated knowledge units $\mathcal{U}_{d}$.
@@ B.1. Datasets
We use two AIS datasets: AIS-DK from the Danish Maritime Authority and AIS-US from the National Oceanic and Atmospheric Administration. AIS-DK covers Danish waters in March 2024, including major routes in the Baltic and North Seas; it contains 10,000 vessel sequences and 2,000,000 records, with an average sequence duration of 0.5 hours across 348 vessels. AIS-US covers U.S. coastal waters in April 2024 with dense traffic near major ports; it contains 10,000 vessel sequences and 2,000,000 records, with an average sequence duration of 2.8 hours across 4,723 vessels. Both datasets include diverse vessel types (e.g., Cargo, Tanker, Passenger), providing representative coverage for evaluation. To construct these datasets, we first downloaded the raw AIS data from the official sources and performed data cleaning to remove abnormal or incomplete records. We then partitioned each vessel trajectory—March 2024 for AIS-DK and April 2024 for AIS-US—into fixed-length segments of 200 records, and uniformly sampled 10,000 segments to build the evaluation corpus.
@@ B.2. Hyperparameters and Implementations
Lin-ITP, Akima Spline, and Kalman Filter are implemented in Python using pandas and pykalman. MH-GIN (Liu et al., 2025a), Multi-task AIS (Nguyen et al., 2018), and KAMEL (Musleh and Mokbel, 2023) follow the authors’ default configurations. GLM-4.5, GLM-4.5-air, Qwen-Plus (snapshot 0112), and Qwen-flash (snapshot 2025-07-28) are accessed through the Aliyun Bailian API (Aliyun Teams, 2025) with thinking mode enabled.
@@ B.3. Evaluation Metrics
We use the Haversine geodesic distance as the joint spatial error and report it in kilometers. The formulas are:
@@ B.4. Baseline Methods
We compare VISTA with three classes of baselines: rule-based trajectory imputation, deep learning-based trajectory imputation, and LLM-based trajectory imputation.
@@
For rule-based trajectory imputation, we include Lin-ITP (Widyantara et al., 2023), Akima Spline (Zaman et al., 2023), and Kalman Filter (Wang et al., 2023). 1) Lin-ITP linearly interpolates longitude and latitude between boundary observations, implicitly assuming constant velocity and heading, and is simple and fast. 2) Akima Spline constructs a piecewise cubic polynomial using local slopes estimated from neighboring points, avoiding oscillations and overshooting common in standard cubic splines while maintaining smooth curvature continuity. It performs better on irregular or nonuniform motion patterns. 3) Kalman Filter treats motion as a linear Gaussian state space with constant velocity dynamics; forward filtering with backward smoothing recovers latent states and positions and is most effective when kinematics are near-linear under moderate Gaussian noise.
@@
For deep learning-based trajectory imputation, we include MH-GIN (Liu et al., 2025a) and Multi-task AIS (Nguyen et al., 2018). MH-GIN is a multi-scale heterogeneous graph imputation network for AIS streams: it extracts multi-scale temporal features per attribute while preserving heterogeneous update rates, then constructs a multi-scale heterogeneous graph to capture cross-attribute dependencies and propagates information to fill missing values. Multi-task AIS is a recurrent framework with latent variable modeling and an embedding of AIS messages designed for high-volume, noisy, and irregularly sampled data; it jointly supports trajectory reconstruction, anomaly detection, and vessel type identification.
@@
For LLM-based trajectory imputation, we include KAMEL (Musleh and Mokbel, 2023), Qwen-plus-th, Qwen-flash-th, GLM-4.5-th, and GLM-4.5-air-th. KAMEL casts trajectory completion as a missing-word problem with spatially aware tokenization and multi-point masked infilling. Beyond KAMEL, to probe the out-of-the-box capability of general LLMs on trajectory imputation without any task-specific design, we adopt two state-of-the-art families, Qwen (Qwen Team, 2025) and GLM 4.5 (Zeng et al., 2025). For each family we include a lightweight and a full-capacity variant, and we enable thinking mode (Alibaba Cloud Team, 2025) throughout (denoted by the “-th” suffix). This yields two paired comparisons under a unified setup: Qwen-plus-th vs Qwen-flash-th for capacity scaling, and GLM-4.5-air-th vs GLM-4.5-th for the trade-off between latency and accuracy.
@@
B.5. Prompt of General Large Language Model for Vessel Trajectory Imputation
@@
Figure 13 presents the baseline prompt used to elicit intermediate points for a missing trajectory segment from a general-purpose LLM. The prompt conditions the model on two boundary observations, previous point ${p}^{-}=[\lambda^{-},\phi^{-}]$ ({prev_str}) and next (first post-gap) point ${p}^{+}=[\lambda^{+},\phi^{+}]$ ({next_str}), and a target count $m$ of missing samples ({missing_length}). The model must output a strict JSON array of $m$ coordinates $\big([\lambda_{1},\phi_{1}],\ldots,[\lambda_{m},\phi_{m}]\big)$, each in [longitude, latitude] with six decimal places, without accompanying text. The constraints emphasize (i) smooth and continuous spatial progression consistent with vessel motion heuristics and (ii) temporal continuity for plausible interpolation between ${p}^{-}$ and ${p}^{+}$. The example block in Figure 13 specifies the exact output schema to ensure deterministic parsing and fair, method-agnostic evaluation of baseline LLMs.
Appendix C Case Study
Figure 14. Case study of VISTA.